find a number such that when 10 is subtracted from twice the number, the result is 14 less than thrice the number.
Answer by harpazo(655)


Let x be the number.
Ten is subtracted from TWICE the number.
2x - 10.
When you do that, the result is 14 less than 3x.

Here is your equation:

2x - 10 = 3x - 14
2x - 3x = -14 + 10
-x = -4
x = -4/-1
x = 4.
The number is 4.